# Semi-Bandit Learning for Monotone Stochastic Optimization

### v1

@@ 1.2 Results and Techniques
The above stability bound can be interpreted as follows: the contribution of item $i$ in the total regret is the probability that it is probed times the total variation error in estimating $\mathcal{D}_{i}$. The dependence on the probability of probing is crucial here. This allows us to analyze the regret using a charging argument where we charge item $i$ for regret in a pay-per-use manner. The paths in the execution tree of the online algorithm that probe $i$ will pay for the error in estimation of $\mathcal{D}_{i}$, but then $\epsilon_{i}$ will also reduce along these paths. Moreover, we can bound the overall regret irrespective of the policy $\sigma$ that is used at periods. The error $\epsilon_{i}\approx\sqrt{1/m}$ when $i$ has been probed $m$ times in the past, which gives a total regret contribution of $\sum_{m=1}^{T}\sqrt{1/m}=O(\sqrt{T})$ for any item. The details appear in §2.1. Finally, we lose an extra logarithmic factor to ensure that our empirical distribution $\mathbf{E}$ stochastically dominates $\mathbf{D}$ with high probability.
@@
Let us contrast our algorithm (and stability bound) with another algorithm that balances the explore-exploit trade-off in a different manner: the explore-then-commit algorithm. This algorithm will explore each item $O(T^{2/3})$ times (say, by probing it first in policy $\sigma$) and then commit to a policy based on these $O(nT^{2/3})$ observations. Using the sample complexity bound of [GHTZ21] it is easy to show that this algorithm achieves $O(T^{2/3})$ regret; see §A for details. However, each item is explored an equal number of times regardless of its cost and we might end up exploring some high-cost spurious items too many times. Our stability bound highlights a crucial difference between our algorithm and this algorithm: we only explore items that are needed in the policy. If there are items which are rarely used in the policy, they will only be explored infrequently and will not contribute to the regret as well. This also explains why our approach achieves a near-optimal $\sqrt{T\log T}$ regret bound compared to the $T^{2/3}$ bound of the sampling-based approach.
@@ 1.3 Related Work
As mentioned earlier, stochastic optimization problems (under known distributions) have been studied extensively. Several papers have extended the classic prophet inequality [SC84] and Pandora’s box [Wei79] results to more complex settings, e.g., [KW12, RS17, KWW16, Sin18, FSZ21, BFLL23]. Moreover, good approximation bounds have been achieved for stochastic versions of various combinatorial optimization problems [DGV08, GK17, Ma18, GKNR15, INZ16, GV06, GN13, GNS17, DHK16, ${}^{+}$, ${}^{+}$12].
@@
There has also been extensive work in online learning, where one considers unknown distributions (in the stochastic learning setting), see e.g. books [CBL06, LS20, Haz22]. All our stochastic problems can be modeled as Markov decision processes (MDPs) with unknown transition probabilities, and there have been some works on achieving sublinear regret in this setting [AOM17, JABJ18, ${}^{+}$20]. However, the regret bounds from these work have a polynomial dependence on the “state space” of the MDP, which in our setting, is exponential in $n$ (the number of r.v.s). It is also known that any online algorithm for arbitrarty MDPs incurs an $\Omega(\sqrt{S\cdot T})$ regret, where $S$ is the size of the state space [JOA10, OR16]. In contrast, we obtain $poly(n)\cdot\sqrt{T\log T}$ regret.
@@
Although both stochastic optimization and online learning have been subjects of comprehensive research, there has been relatively limited work at the intersection of these two domains. For instance, [${}^{+}$13] studied a special case of adaptive submodular maximization [GK17] in the bandit setting. Some recent papers [${}^{+}$20, GT22] considered online learning for Pandora’s box in a correlated setting, which is much harder than the independent case: one needs to resort to weaker “partially adaptive” benchmarks here. [EHLM19] also considered a different “multi armed” Pandora’s box problem, and obtained a $\frac{1}{2}$-competitive ratio relative to a suitable benchmark. Drawing closer to our work is [GKSW24], which explores Pandora’s box and prophet inequality within the online learning framework with bandit feedback, which is even more restrictive than our semi-bandit feedback. In bandit feedback, the algorithm only observes the realized objective value of its policy $\sigma$. [GKSW24] gave algorithms with $1$-regret $O\left(n^{3}\sqrt{T}\log(T)\right)$ and $O\left(n^{4.5}\sqrt{T}\log(T)\right)$ for Pandora’s box and prophet inequality respectively. While our feedback model is not as general as [GKSW24], we think that semi-bandit feedback already captures the issue of partial observations in policies. Moreover, under semi-bandit feedback we obtain a general framework for online learning algorithms, that applies to several stochastic optimization problems.
@@
There have also been several works studying sample complexity bounds for several stochastic optimization problems. The goal in these works is to understand how much data is necessary and sufficient to guarantee near-optimal algorithms. Such results have been obtained for single-parameter revenue maximization [CR14, RS16] and prophet inequality [CDFS19]. Recent work [GHTZ21] gives optimal sample complexity bounds for stochastic optimization problems that exhibit strong monotonicity, which is a slightly stronger condition than our monotonicity definition. These results imply $\sqrt{T}$ regret under the full-feedback model, but only ${T}^{2/3}$ regret under the semi-bandit feedback model considered in this paper (see Appendix A).
@@
Our results are related in spirit to some previous online-to-offline reductions. In particular, [KV05] considered combinatorial optimization with linear objective functions, and obtained $1$-regret of $O(n\sqrt{T})$ under full-feedback (assuming an exact offline algorithm). Then, [KKL09] extended this result to linear problems with only an $\alpha$-approximation algorithm, and obtained $\alpha$-regret of $O(n\sqrt{T})$ under full-feedback and $O(nT^{2/3})$ under bandit-feedback. Recent work of [${}^{+}$21] considers certain combinatorial optimization problems with non-linear objectives, assuming an $\alpha$-approximation via a greedy-type algorithm. For such problems, [${}^{+}$21] obtained $\sqrt{T}$ regret under full-feedback and $T^{2/3}$ regret under bandit-feedback. We note that all these results work in the (harder) adversarial online setting, rather than our i.i.d. online setting. However, our result handles a much wider class of problems, with arbitrary objectives. Moreover, as noted before, there is an $\Omega(T)$ regret lower-bound for many of our applications in the adversarial setting.
@@ 1.4 Preliminaries
A probability distribution $\mathcal{E}$ (over $\mathbb{R}$) stochastically dominates another distribution $\mathcal{D}$ if, for all $a\in\mathbb{R}$, we have: $\mathbf{P}_{X\leftarrow\mathcal{E}}(X\geq a)\geq\mathbf{P}_{Y\leftarrow\mathcal{D}}(Y\geq a)$.
We use $\mathcal{D}\preceq_{\mathsf{SD}}\mathcal{E}$ to denote that distribution $\mathcal{E}$ stochastically dominates $\mathcal{D}$. For product distributions $\mathbf{{D}}=\{\mathcal{D}_{i}\}_{i=1}^{n}$ and $\mathbf{{E}}=\{\mathcal{E}_{i}\}_{i=1}^{n}$, we say that $\mathbf{{E}}$ stochastically dominates $\mathbf{{D}}$ if $\mathcal{E}_{i}$ stochastically dominates $\mathcal{D}_{i}$ for all $i\in[n]$; we also denote this by $\mathbf{{D}}\preceq_{\mathsf{SD}}\mathbf{{E}}$.
Probability Distribution Metric.
Let $\mathtt{TV}(\mathcal{D},\mathcal{E})$ denote the total variation distance between discrete distributions $\mathcal{D}$ and $\mathcal{E}$. The total variation distance is half of the $\ell_{1}$ distance between the two distributions, i.e., $\mathtt{TV}(\mathcal{D},\mathcal{E})=\frac{1}{2}\cdot||\mathcal{D}-\mathcal{E}||_{1}$. The following standard result (see, for example, Lemma B.8 in[GVdV17]) bounds the total variation distance between product distributions.
@@
$\mathtt{TV}(\mathbf{{D}},\mathbf{{E}})\leq\sum_{i\in[n]}\mathtt{TV}(\mathcal{D}_{i},\mathcal{E}_{i}).$
Consider independent r.v.s $\mathsf{X}_{1},\ldots,\mathsf{X}_{n}$ where $\mathsf{T}_{i}$ denotes the domain of $\mathsf{X}_{i}$. Let $h$ be a function from $\mathbf{{T}}=\mathsf{T}_{1}\times\cdots\times\mathsf{T}_{n}$ to $[0,U]$; that is, $h$ is a function on the outcomes of the random variables that is bounded by $U$. Thus, for any $\mathbf{x}\in\mathbf{{T}}$, $h(\mathbf{x})$ denotes the value of $h$ on the outcome $\mathbf{x}=(x_{1},\ldots x_{n})$. Given a product distribution $\mathbf{{P}}$ over the r.v.s, define
$h(\mathbf{{P}})=\mathbb{E}_{\mathbf{x}\sim\mathbf{{P}}}\left[h(\mathbf{x})\right].$
The following useful fact bounds the difference in function value at two different distributions.
@@
Given discrete distributions $\mathbf{{D}}$ and $\mathbf{{E}}$ over $n$ random variables $\mathsf{X}_{1},\ldots,\mathsf{X}_{n}$, we have
$|h(\mathbf{{D}})-h(\mathbf{{E}})|\,\,\leq\,\,U\cdot\mathtt{TV}(\mathbf{{D}},\mathbf{{E}}).$
@@ 2 The Online Framework
There is an algorithm $\mathtt{EmpStocDom}$ that, given $m$ i.i.d. samples from a distribution $\mathcal{D}$ with finite support-size $k$, and parameter $\delta>0$, computes a distribution $\mathcal{E}$ that satisfies the following properties with probability at least $1-\delta$:
@@
  • the total-variation distance $\mathtt{TV}(\mathcal{E},\mathcal{D})<k\sqrt{\frac{\log(2k/\delta)}{2m}}$.
@@
We will assume, without loss of generality, that we have observed at least one sample from each random variable. This can be ensured by adding $n$ special time periods and choosing for each time $i\in[n]$, a policy $\sigma^{i}$ that first probes item $i$; note that this contributes at most $n\cdot f_{max}$ to the total regret. Let $N_{j}^{t}$ be the number of times r.v. $\mathsf{X}_{j}$ has been sampled before time $t$. By our assumption, $N_{j}^{1}=1$ for all $j\in[n]$. Note that $N_{j}^{t}$ is a random variable and depends on the history $\mathbf{h}^{t-1}=\left(\mathbf{x}^{1},\cdots\mathbf{x}^{t-1}\right)$. When needed, we will use $N_{j}^{t}(\mathbf{h}^{t-1})$ to make this dependence explicit. The following result uses Theorem 2.1 and union bound.
@@
$\mathcal{E}_{j}^{t}\text{ stochastically dominates }\mathcal{D}_{j}\text{ and }\mathtt{TV}\left(\mathcal{E}^{t}_{j},\mathcal{D}_{j}\right)<k\cdot\sqrt{\frac{3\log(knT)}{2N_{j}^{t}}},\,\,\mbox{ for all }j\in[n]\mbox{ and }t\in[T].$
@@
$\mathbf{P}\left(B_{j}^{t}\ \wedge\ N_{j}^{t}=m\right)\leq\delta,\qquad\forall m\geq 1.$
@@
$\mathbf{P}(B_{j}^{t})\leq\sum_{m=1}^{T}\mathbf{P}\left(B_{j}^{t}\ \wedge\ N_{j}^{t}=m\right)\leq\sum_{m=1}^{T}\delta=\delta T.$
@@
$\mathbf{P}\left(\vee_{j=1}^{n}\vee_{t=1}^{T}B_{j}^{t}\right)\leq\sum_{j=1}^{n}\sum_{t=1}^{T}\mathbf{P}(B_{j}^{t})\leq nT^{2}\delta\leq\frac{1}{nT}.$
@@
We now define a good event $G$ that corresponds to the condition in Lemma 2.2 holding true (for all $j\in[n]$ and $t\in[T]$). Next, we complete the proof assuming that $G$ holds. At the end, we show how to remove this assumption (using the fact that $G$ holds with high probability).
@@
Consider a stochastic problem that is up-monotone. Suppose that $\mathbf{E}=\{\mathcal{E}_{i}\}_{i=1}^{n}$ and $\mathbf{D}=\{\mathcal{D}_{i}\}_{i=1}^{n}$ are product distributions such that $\mathbf{D}\preceq_{\mathsf{SD}}\mathbf{E}$ and $\mathtt{TV}(\mathcal{E}_{i},\mathcal{D}_{i})\leq\epsilon_{i}$ for each $i\in[n]$. If $\sigma$ is the policy returned by $\mathtt{ALG}(\mathbf{E})$ and $\sigma^{*}$ is an optimal policy under $\mathbf{D}$, then:
@@
This lemma relies on the monotonicity assumption, and is proved in §2.2.
2.1 Completing the Proof of Theorem 1.1
We now complete the proof of Theorem 1.1 using Lemma 2.3.
Bounding the regret as sum over time $t$.
@@
where $\mathbf{h}^{t-1}$ is the history at time $t$, policy $\sigma^{t}=\mathtt{ALG}(\mathbf{E}^{t})$ and policy $\sigma^{*}$ is the optimal policy for the stochastic problem instance (under the true distribution $\mathbf{D}$), and policy $\sigma^{t}$ is the policy used by our online algorithm at time $t$. By (3), the overall regret is
$\alpha\text{-}R(T)=\sum_{t=1}^{T}\mathbb{E}_{\mathbf{h}^{t-1}}\left[R^{t}(\mathbf{h}^{t-1})\right].$
For each time $t\in[T]$ and history $\mathbf{h}^{t-1}$, we apply Lemma 2.3 with distributions $\mathbf{E}^{t}(\mathbf{h}^{t-1})$ and $\mathbf{D}$, and parameters $\epsilon_{i}^{t}(\mathbf{h}^{t-1})=k\cdot\sqrt{\frac{3\log(knT)}{2N_{i}^{t}(\mathbf{h}^{t-1})}}$ for all $i\in[n]$. Note that $\mathtt{TV}(\mathcal{E}^{t}_{i},\mathcal{D}_{i})\leq\epsilon_{i}^{t}(\mathbf{h}^{t-1})$ and $\mathcal{D}_{i}\preceq_{\mathsf{SD}}\mathcal{E}_{i}$ for each $i\in[n]$ because we assumed the good event $G$. Hence, Lemma 2.3 implies that:
@@
We note that $R^{t}$, $\sigma^{t}$, $N^{t}$ and $\epsilon^{t}$ all depend on the history $\mathbf{h}^{t-1}$; to keep notation simple, we drop the explicit dependence on $\mathbf{h}^{t-1}$. Combining the above two inequalities,
$\alpha\text{-}R(T)\leq f_{\max}\sum_{t=1}^{T}\mathbb{E}_{\mathbf{h}^{t-1}}\left[\sum_{i=1}^{n}{Q}_{i}(\sigma^{t})\cdot\epsilon_{i}^{t}\right]=kf_{\max}\sqrt{1.5\log(knT)}\cdot\sum_{t=1}^{T}\mathbb{E}_{\mathbf{h}^{t-1}}\left[\sum_{i=1}^{n}\frac{{Q}_{i}(\sigma^{t})}{\sqrt{N^{t}_{i}}}\right].$
@@
Indeed, combining (6) with the above bound on regret, we get $\alpha\text{-}R(T)\leq knf_{max}\sqrt{6\log(knT)}$, which completes the proof of Theorem 1.1.
Proving (6) as sum over decision paths.
@@
$=\sum_{t=1}^{T}\mathbb{E}_{\mathbf{h}^{t-1}}\left[\frac{1}{\sqrt{N^{t}_{i}}}\cdot\mathbf{P}_{\mathbf{x}^{t}}[i\text{ probed by }\sigma^{t}]\right]\,\,=\,\,\sum_{t=1}^{T}\mathbb{E}_{\mathbf{h}^{t-1}}\left[\frac{Q_{i}(\sigma^{t})}{\sqrt{N^{t}_{i}}}\right],$ (8)
@@
The first inequality uses the fact that $N^{t}_{i}(\mathbf{h}^{t-1})$ equals the number of probes of item $i$ in the first $t-1$ time steps: so $N^{t+1}_{i}=N^{t}_{i}+1$ whenever item $i$ is probed by $\sigma^{t}(\mathbf{h}^{t-1})$. This completes the proof of (9) and hence (6).

### v2

@@ 1.2 Main Result and Techniques
The above stability bound can be interpreted as follows: the contribution of item $i$ in the total regret is the probability that it is probed times the total variation error in estimating $\mathcal{D}_{i}$. The dependence on the probability of probing is crucial here. This allows us to analyze the regret using a charging argument where we charge item $i$ for regret in a pay-per-use manner. The paths in the execution tree of the online algorithm that probe $i$ will pay for the error in estimation of $\mathcal{D}_{i}$, but then $\epsilon_{i}$ will also reduce along these paths. Moreover, we can bound the overall regret irrespective of the policy $\sigma$ that is used at periods. The error $\epsilon_{i}\approx\sqrt{1/m}$ when $i$ has been probed $m$ times in the past, which gives a total regret contribution of $\sum_{m=1}^{T}\sqrt{1/m}=O(\sqrt{T})$ for any item. Finally, we lose an extra logarithmic factor to ensure that our empirical distribution $\mathbf{E}$ stochastically dominates $\mathbf{D}$ with high probability.
@@
We contrast our algorithm (and stability bound) with another algorithm that balances the explore-exploit trade-off in a different manner: the explore-then-commit algorithm. Using the sample complexity bound of Guo et al. (2021) it is easy to show that this algorithm achieves $T^{2/3}$ regret; see Appendix A for details. However, each item is explored an equal number of times regardless of its quality and we might end up exploring some “low quality” items too many times. Our stability bound highlights a crucial difference between our algorithm and explore-then-commit: we only explore items that are actually probed by the (current) optimal policy. If there are low-quality items that are rarely used in the optimal policy, they will be explored infrequently in our algorithm: so their regret contribution is low. This explains why our approach achieves a near-optimal $\sqrt{T\log T}$ regret bound compared to the $T^{2/3}$ bound of the sampling-based approach.
1.3 Additional Results and Applications
We can also extend our main result to stochastic problems with continuous (or mixed) distributions. Here, we need to make some additional (mild) assumptions on the objective function. Informally, we require that for any fixed policy and leaf-node, the function value does not vary significantly with the variables $\mathbf{x}$. This assumption is satisfied whenever the function value at a leaf-node is linear or monotone, which suffices for applications such as Prophet inequality and Pandora’s box. In this case, we are able to obtain $\alpha$-regret bounds similar to Theorem 1.1. The algorithm for the continuous setting is identical to the discrete case. The main new idea in the analysis is the use of a different distance-measure between distributions (the Kolmogorov–Smirnov distance) and modifying the proof of the “stability” bound appropriately. We defer the formal definitions and results to Section 3.
Another extension of our main result is to censored and binary feedback. This extension requires discrete distributions and a more restrictive assumption on the objective function. Roughly speaking, we assume that the function value can always be determined based on the observed feedback (either censored or binary). This assumption is satisfied for the single-resource revenue management and sequential posted pricing problems. Again, we obtain regret bounds as in Theorem 1.1. While the high-level algorithm is the same, we need to design new sampling algorithms to construct the stochastically dominating distribution $\mathbf{{E}}$. This requires more work because the feedback from a probed item $i$ does not directly provide us with a sample of $\mathsf{X}_{i}$. All the details of this extension are in Section 4. We note that the discrete-distribution assumption is necessary here in order to obtain $\sqrt{T}$ regret: see Appendix B for a regret lower bound of $\Omega(T^{2/3})$ under binary-feedback and continuous distributions.
We demonstrate the generality of our framework by obtaining polynomial time online learning algorithms for a number of stochastic optimization problems arising in domains of optimal stopping, revenue management, submodular optimization, knapsack and matching. For many problems, we get better regret bounds using our framework than what was known previously using problem-specific methods. We summarize our results in Table 1 and refer to Section 5 for further details.
@@ 1.4 Related Work
As mentioned earlier, stochastic optimization problems (under known distributions) have been studied extensively. Several papers have extended the classic prophet inequality (Samuel-Cahn, 1984) and Pandora’s box (Weitzman, 1979) results to more complex settings, e.g., Kleinberg and Weinberg (2012); Rubinstein and Singla (2017); Kleinberg et al. (2016); Singla (2018); Feldman et al. (2021); Boodaghians et al. (2023). Moreover, good approximation bounds have been achieved for stochastic versions of various combinatorial optimization problems (Dean et al., 2008; Golovin and Krause, 2017; Ma, 2018; Gupta et al., 2015; Im et al., 2016; Goemans and Vondrák, 2006; Gupta and Nagarajan, 2013; Gupta et al., 2017; Deshpande et al., 2016; Chen et al., 2009; Bansal et al., 2012).
@@
There has also been extensive work in online learning, where one considers unknown distributions (in the stochastic learning setting), see e.g., books Cesa-Bianchi and Lugosi (2006), Lattimore and Szepesvári (2020) and Hazan (2022). All our stochastic problems can be modeled as Markov decision processes (MDPs) with unknown transition probabilities, and there have been some works on achieving sublinear regret in this setting (Azar et al., 2017; Jin et al., 2018, 2020). However, the regret bounds from these works have a polynomial dependence on the “state space” of the MDP, which in our setting, is exponential in $n$ (the number of r.v.s). It is also known that any online algorithm for arbitrary MDPs incurs an $\Omega(\sqrt{S\cdot T})$ regret, where $S=\exp(n)$ is the size of the state space (Jaksch et al., 2010; Osband and Roy, 2016). In contrast, we obtain regret bounds that depend polynomially on $n$.
@@
Although both stochastic optimization and online learning have been subjects of comprehensive research, there has been relatively limited work at the intersection of these two domains. Gatmiry et al. (2024) obtained online learning algorithms for Pandora’s box and prophet inequality under bandit feedback, which is even more restrictive than semi-bandit feedback. Here, the algorithm only observes the realized objective value of its policy. Gatmiry et al. (2024) gave algorithms with $1$-regret $O\left(n^{3}\sqrt{T}\log(T)\right)$ and $O\left(n^{4.5}\sqrt{T}\log(T)\right)$ for Pandora’s box and prophet inequality respectively. While our semi-bandit feedback model is more relaxed than bandit feedback, we think that it already captures the core issue of partial observations in learning policies. Moreover, under semi-bandit feedback we are able to obtain a general framework that applies to several stochastic optimization problems.
@@
There have also been several works studying sample complexity bounds for stochastic optimization problems. The goal in these works is to understand how much data is necessary and sufficient to guarantee near-optimal algorithms. Such results have been obtained for single-parameter revenue maximization (Cole and Roughgarden, 2014; Roughgarden and Schrijvers, 2016) and prophet inequality (Correa et al., 2019). The recent work of Guo et al. (2021) gives optimal sample complexity bounds for stochastic optimization problems that exhibit strong monotonicity, which is a slightly stronger condition than our monotonicity definition.
Another relevant line of work is on combinatorial multi-armed bandits (CMAB), which also involves semi-bandit feedback (Kveton et al., 2014, 2015; Chen et al., 2016b, a). Here, there are $n$ base arms that produce stochastic outcomes drawn from an unknown fixed distribution. There is also a collection ${\cal F}\subseteq 2^{[n]}$ of allowed “super arms”. In each period $t$, the algorithm selects a super-arm $S^{t}\in{\cal F}$ and observes the realizations of all arms $i\in S^{t}$. The result closest to ours is by Chen et al. (2016a), which considers a class of non-linear objectives satisfying a “monotone” condition and obtains a UCB-type algorithm achieving $\alpha$-regret of $O(n\sqrt{T\log T})$ where $\alpha$ is the approximation ratio for the offline problem. Our setting is more general because we need to select policies (not static subsets). When we select a policy, we do not know which arms will actually be observed. So, we only have an indirect control on what arms will be observed in each round. While our algorithm can be seen as a natural extension of Chen et al. (2016a), our analysis requires new ideas: in particular in proving the stability Lemma 2.3 and its use in bounding overall regret.
@@
There are also some online-to-offline reductions that work in the adversarial setting (which is harder than our stochastic setting). In particular, Kalai and Vempala (2005) considered combinatorial optimization with linear objective functions, and obtained $1$-regret of $O(n\sqrt{T})$ under full-feedback (assuming an exact offline algorithm). Then, Kakade et al. (2009) extended this result to linear problems with only an $\alpha$-approximation algorithm, and obtained $\alpha$-regret of $O(n\sqrt{T})$ under full-feedback and $O(nT^{2/3})$ under bandit-feedback. Recent work of Niazadeh et al. (2021); Nie et al. (2023) considers certain combinatorial optimization problems with non-linear objectives, assuming an $\alpha$-approximation via a greedy-type algorithm. For such problems, Niazadeh et al. (2021) obtained $\sqrt{T}$ regret under full-feedback and $T^{2/3}$ regret under bandit-feedback. While these results work in the harder adversarial online setting, our result handles a much wider class of problems: we learn policies (rather than just subsets) and handle arbitrary objectives. As noted before, there is an $\Omega(T)$ regret lower-bound for some of our applications in the adversarial setting.
@@ 1.5 Preliminaries
A probability distribution $\mathcal{E}$ (over $\mathbb{R}$) stochastically dominates another distribution $\mathcal{D}$ if, for all $a\in\mathbb{R}$, we have: $\Pr_{\mathsf{X}\sim\mathcal{E}}(\mathsf{X}\geq a)\geq\Pr_{\mathsf{Y}\sim\mathcal{D}}(\mathsf{Y}\geq a)$.
[…]
@@
$\operatorname{\mathtt{TV}-\mathtt{Dist}}(\mathbf{{D}},\mathbf{{E}})\leq\sum_{i\in[n]}\operatorname{\mathtt{TV}-\mathtt{Dist}}(\mathcal{D}_{i},\mathcal{E}_{i}).$
Consider independent r.v.s $\mathsf{X}_{1},\ldots,\mathsf{X}_{n}$ where $\mathsf{T}_{i}$ denotes the domain of $\mathsf{X}_{i}$. Let $h$ be a function from $\mathbf{{T}}=\mathsf{T}_{1}\times\cdots\times\mathsf{T}_{n}$ to $[0,U]$; that is, $h$ is a function on the outcomes of the random variables that is bounded by $U$. Thus, for any $\mathbf{x}\in\mathbf{{T}}$, $h(\mathbf{x})$ denotes the value of $h$ on the outcome $\mathbf{x}=(x_{1},\ldots x_{n})$. Given a product distribution $\mathbf{{P}}$ over the r.v.s, define $h(\mathbf{{P}}):=\mathbb{E}_{\mathbf{x}\sim\mathbf{{P}}}\left[h(\mathbf{x})\right]$. The following useful fact bounds the difference in function value at two different distributions.
@@
Given discrete distributions $\mathbf{{D}}$ and $\mathbf{{E}}$ over $n$ random variables $\mathsf{X}_{1},\ldots,\mathsf{X}_{n}$, and a $[0,U]$ bounded function $h$ on the outcomes of these r.v.s (as above), we have
$|h(\mathbf{{D}})-h(\mathbf{{E}})|\,\,\leq\,\,U\cdot\operatorname{\mathtt{TV}-\mathtt{Dist}}(\mathbf{{D}},\mathbf{{E}}).$
Theorem 1.4 (DKW Inequality from Massart (1990)).
Let $X_{1},\ldots,X_{N}$ be i.i.d. samples with cumulative distribution function $F(\cdot)$. Let $\overline{F}(x):=\frac{1}{N}\sum_{i\in[N]}\mathbf{1}_{X_{i}\leq x}$ be the empirical CDF. Then, for every $\varepsilon>0$, we have
$\Pr\left[\sup_{x}\left|\overline{F}(x)-F(x)\right|\,>\,\varepsilon\right]\,\,\leq\,\,2\exp\left(-2N\varepsilon^{2}\right)$
@@ 2 The Online Framework
There is an algorithm $\mathtt{EmpStocDom}$ that, given $m$ i.i.d. samples from a distribution $\mathcal{D}$ with finite support-size $k$, and parameter $\delta>0$, computes a distribution $\mathcal{E}$ that satisfies the following properties with probability at least $1-\delta$.
@@
  • the total-variation distance $\operatorname{\mathtt{TV}-\mathtt{Dist}}(\mathcal{E},\mathcal{D})<k\sqrt{\frac{\log(2/\delta)}{2m}}$.
@@
Proof of Theorem 1.1.
We will assume, without loss of generality, that we have observed at least one sample from each random variable. This can be ensured by adding $n$ special time periods and choosing for each time $i\in[n]$, a policy $\sigma^{i}$ that first probes item $i$; note that this contributes at most $n\cdot f_{max}$ to the total regret. Let $N_{j}^{t}$ be the number of times r.v. $\mathsf{X}_{j}$ has been sampled before time $t$. By our assumption, $N_{j}^{1}=1$ for all $j\in[n]$. Note that $N_{j}^{t}$ is a random variable and depends on the history $\mathbf{h}^{t-1}=\left(\mathbf{x}^{1},\cdots\mathbf{x}^{t-1}\right)$. When needed, we will use $N_{j}^{t}(\mathbf{h}^{t-1})$ to make this dependence explicit. The following result follows directly from Theorem 2.1.
@@
[…]
@@
$\Pr\left(B_{j}^{t}\ \wedge\ N_{j}^{t}=m\right)\leq\delta,\qquad\forall m\geq 1.$
@@
$\Pr(B_{j}^{t})\leq\sum_{m=1}^{T}\Pr\left(B_{j}^{t}\ \wedge\ N_{j}^{t}=m\right)\leq\sum_{m=1}^{T}\delta=\delta T.$
@@
$\Pr\left(\vee_{j=1}^{n}\vee_{t=1}^{T}B_{j}^{t}\right)\leq\sum_{j=1}^{n}\sum_{t=1}^{T}\Pr(B_{j}^{t})\leq nT^{2}\delta\leq\frac{1}{nT}.$
@@
Let $G$ denote the “good” event corresponding to the condition in Lemma 2.2 holding for all $j$ and $t$. First, we complete the proof assuming that $G$ holds (this assumption is removed later).
@@
Consider a stochastic problem that is up-monotone. Suppose that $\mathbf{E}=\{\mathcal{E}_{i}\}_{i=1}^{n}$ and $\mathbf{D}=\{\mathcal{D}_{i}\}_{i=1}^{n}$ are product distributions such that $\mathbf{D}\preceq_{\mathsf{SD}}\mathbf{E}$ and $\operatorname{\mathtt{TV}-\mathtt{Dist}}(\mathcal{E}_{i},\mathcal{D}_{i})\leq\epsilon_{i}$ for each $i\in[n]$. If $\sigma$ is the policy returned by $\mathtt{ALG}(\mathbf{E})$ and $\sigma^{*}$ is an optimal policy under $\mathbf{D}$, then:
@@
This lemma relies on the monotonicity assumption, and is proved in §2.1. We now complete the proof of Theorem 1.1 using Lemma 2.3.
Bounding the regret as a sum over time $t$.
@@
where $\mathbf{h}^{t-1}$ is the history at time $t$, policy $\sigma^{t}=\mathtt{ALG}(\mathbf{E}^{t})$ and policy $\sigma^{*}$ is the optimal policy for the stochastic problem instance (under the true distribution $\mathbf{D}$). By (3), the overall regret is $\alpha\text{-}R(T)=\sum_{t=1}^{T}\mathbb{E}_{\mathbf{h}^{t-1}}\left[R^{t}(\mathbf{h}^{t-1})\right]$. For each time $t\in[T]$ and history $\mathbf{h}^{t-1}$, we apply Lemma 2.3 with distributions $\mathbf{E}^{t}(\mathbf{h}^{t-1})$ and $\mathbf{D}$, and parameters $\epsilon_{i}^{t}(\mathbf{h}^{t-1})=k\cdot\sqrt{\frac{3\log(nT)}{2N_{i}^{t}(\mathbf{h}^{t-1})}}$ for all $i\in[n]$. Note that $\operatorname{\mathtt{TV}-\mathtt{Dist}}(\mathcal{E}^{t}_{i},\mathcal{D}_{i})\leq\epsilon_{i}^{t}(\mathbf{h}^{t-1})$ and $\mathcal{D}_{i}\preceq_{\mathsf{SD}}\mathcal{E}_{i}$ for each $i\in[n]$ because we assumed the good event $G$. Hence, Lemma 2.3 implies that:
@@
We note that $R^{t}$, $\sigma^{t}$, $N^{t}$ and $\epsilon^{t}$ all depend on the history $\mathbf{h}^{t-1}$; to keep notation simple, we drop the explicit dependence on $\mathbf{h}^{t-1}$. Using the regret definition and (5),
$\alpha\text{-}R(T)\leq f_{\max}\sum_{t=1}^{T}\mathbb{E}_{\mathbf{h}^{t-1}}\left[\sum_{i=1}^{n}{Q}_{i}(\sigma^{t})\cdot\epsilon_{i}^{t}\right]=kf_{\max}\sqrt{1.5\log(nT)}\cdot\sum_{t=1}^{T}\mathbb{E}_{\mathbf{h}^{t-1}}\left[\sum_{i=1}^{n}\frac{{Q}_{i}(\sigma^{t})}{\sqrt{N^{t}_{i}}}\right].$
@@
Indeed, combining (6) with the above bound on regret, we get $\alpha\text{-}R(T)\leq knf_{max}\sqrt{6T\log(nT)}$, which would complete the proof of Theorem 1.1 (assuming event $G$).
Proving (6) as a sum over decision paths.
@@
$=\sum_{t=1}^{T}\mathbb{E}_{\mathbf{h}^{t-1}}\left[\frac{1}{\sqrt{N^{t}_{i}}}\cdot\Pr_{\mathbf{x}^{t}}[i\text{ probed by }\sigma^{t}]\right]\,\,=\,\,\sum_{t=1}^{T}\mathbb{E}_{\mathbf{h}^{t-1}}\left[\frac{Q_{i}(\sigma^{t})}{\sqrt{N^{t}_{i}}}\right],$ (8)
@@
The first inequality uses the fact that $N^{t}_{i}(\mathbf{h}^{t-1})$ equals the number of probes of item $i$ in the first t1t-1 time steps: so $N^{t+1}_{i}=N^{t}_{i}+1$ whenever item $i$ is probed by $\sigma^{t}(\mathbf{h}^{t-1})$. This completes the proof of (9) and hence (6).